OpenStudy (anonymous):

Simplify the expression. -5+i/2i

OpenStudy (zarkon):

\[\Huge\frac{-5+i}{2i}\] ?

OpenStudy (anonymous):

Yeah. . .

OpenStudy (zarkon):

multiply top and bottom by \(i\)

OpenStudy (zarkon):

then use the fact that \(i^2=-1\)

OpenStudy (anonymous):

-10i?

OpenStudy (anonymous):

Hey. . .I'm not dun with you. : I

OpenStudy (zarkon):

a better way to say that is...I still need more help

OpenStudy (zarkon):

people here provide their expertise free of charge

OpenStudy (zarkon):

where did you get -10i from

OpenStudy (anonymous):

You said multiply top and bottom. . .

OpenStudy (zarkon):

\[\frac{-5+i}{2i}\frac{i}{i}=\frac{-5i+i^2}{2i^2}\]

OpenStudy (zarkon):

then use, like I said above, \(i^2=-1\)

OpenStudy (anonymous):

How do I do that?

OpenStudy (anonymous):

Zaaaaaaaaaaaaaaaaaaaaaaaaaaaaarrrrrrrrrrrrrrrrrraaaaa!

OpenStudy (anonymous):

-5(-1)+(-1) 2(-1) ?

OpenStudy (anonymous):

At least complete the task at hand, if you're going to help. >~>

OpenStudy (zarkon):

if I do anymore I will just be giving you the answer...and I'm not going to do that.

OpenStudy (anonymous):

. . .What a drag.